OpenStudy (anonymous):

10u^3-5u^2=0 Solve pls help

OpenStudy (anonymous):

2 solutions. 10u^3 - 5u^2 = 0 factor out a 5u^2 5u^2 (2u - 1) = 0 5u^2 = 0 and (2u - 1) = 0 u = 0 and 2u = 1 u = 0 and u = (1/2)